OpenStudy (anonymous):

Layla has a jar filled with 108 coins. If the jar contains only quarters and nickels and has a value of $15.80, how many nickels are in the jar 52 53 55 56

OpenStudy (anonymous):

I thought what I was supposed to do was set up 2 equations so I did x+y=108, .5x+.25y=15.80. Then I thought it would be easier just to multiple the answers by .5 then subtract that from 15.80 and figure out how many quarters there would be and make sure it added up to 108. but .5*52=26 which is already more than 15.80. So I am not sure..

OpenStudy (anonymous):

let x be the # of nickels let y be the # of quarters you can form two equations, the first for the total number of coins: \[x +y = 108\]\[x = 108 -y\] The second for the total amount: \[(NumNickels) * (ValNickels) + (NumQuarters)*(ValQuarters) = $15.80\] \[(x)*(0.05)+(y)*($0.25)=$15.80\]

OpenStudy (anonymous):

Oh. I see what I did wrong now. Thank you so much! You are amazing!